if you are an analyst at the ministry of transport and the minister proposes that the toll of 50p would now be 1.00ghc and the elasticity of demand for the motorway is 0.3. Do you think this is politically a good decision?



2. Do you think the price support system would help farmers in Ghana?



3. Explain the concept of market failure.


1
Expert's answer
2021-03-23T11:49:50-0400

Solution:

1.). Price elasticity of demand = "\frac{Percentage\; change\; in \;quantity\; demanded}{Percentage\; change\; in\; price}"


Price elasticity of demand = 0.3

Percentage change in price = "\frac{Change\; in \;price}{Price} = \frac{1 - 50}{50} = 1"


Percentage change in quantity demanded ="\frac{Change\; in \;quantity}{Quantity} = ?"

Therefore:


0.3 = "\frac{Percentage\; change\; in \;quantity\; demanded}{1}"

Percentage change in quantity demanded = 0.3 < 1

This means that demand for motorways is inelastic, that is a change in price causes a smaller percentage change in demand.

Therefore, this is politically a good decision, since the increase in price slightly affected the quantity demanded, which was minimal.

 

2.). Price supports refer to price controls or subsidies that are leveraged by the government to increase or decrease prices and hence alter the supply consumed or quantity demanded by individuals within the economic system.

Therefore, the price support system will help farmers in Ghana by stabilizing prices. The Ghana government through buffer stock operations can protect smallholder farmer’s income. The stable prices will motivate the farmers to produce more since they have a stable market for their produce.

 

3.). Market failure is an economic situation described as an inefficient distribution of goods and services in the free market. It occurs when there is a state of disequilibrium in the market due to market distortion. It manifests when the quantity of goods or services supplied is not equal to the quantity of goods or services demanded.